OpenStudy (anonymous):

System of Equations Equation I: y = 3x + 5 Equation II: x + y = 3 Solve the above system of equations by using the subsitution method. Show all your work.

OpenStudy (anonymous):

substitute value of y in 2º equation : x+3x + 5 =3 and solve for x: 4x=-2 x=-1/2 put this value of x in 1º equation and solve for y: y=3(-1/2)+5=-3/2+10/2=7/2